Please solve it hand written no excel
The refinery has U kerosene, and 650,000 barrel a) Formulate the problem of finding the numbl demand and minimizes the total cost as a linear program b) Solve the problem using the graphic method 2. (25 points) Toolco has contracted with AutoMate to supply their automotive discount stores with wrenches and chisels. AutoMates weekly demand consists of at least 1500 wrenches and 1200 chisels. Toolco cannot produce all requested units with its present one-shift capacity and must use overtime and possibly subcontract with other tool shops. The result is an increase in the production cost per unit, as shown in the following table. Market demand restricts the ratio chisels to wrenches to at least 2:1 (chisels at least 2 times the number of wrenches) Tool Production type Weekly production Unit cost range (units 2.00 2.8 0-550 Regular Over time Subcontractin Regular Over time Subcontracting Wrenches 551-800 801-00 0-620 621-900 3.00 2.10 3.2 4.2 Chisel 901-0o Formulate the problem as a line each tool. ar program to determine the optimum production schedule f
Answer #1

Linear program is following:

Let W1, W2, W3 be the number of Wrenches to be produced in Regular-time, Overtime and Subcontracting respectively

C1, C2, C3 be the number of Chisels to be produced in Regular-time, Overtime and Subcontracting respectively

Objective: Min 2W1+2.8W2+3W3+2.1C1+3.2C2+4.2C3

s.t.

W1+W2+W3 >= 1500

C1+C2+C3 >= 1200

C1+C2+C3 >= 2(W1+W2+W3)  

or, C1+C2+C3-2W1-2W2-2W3 >= 0

W1 <= 550

W2 <= 250 (=800-550)

C1 <= 620

C2 <= 280 (=900-620)

All variables >= 0